# In-context Learning in Presence of Spurious Correlations

### v1

@@ Abstract
Large language models exhibit a remarkable capacity for in-context learning, where they learn to solve tasks given a few examples. Recent work has shown that transformers can be trained to perform simple regression tasks in-context. This work explores the possibility of training an in-context learner for classification tasks involving spurious features. We find that the conventional approach of training in-context learners is susceptible to spurious features. Moreover, when the meta-training dataset includes instances of only one task, the conventional approach leads to task memorization and fails to produce a model that leverages context for predictions. Based on these observations, we propose a novel technique to train such a learner for a given classification task. Remarkably, this in-context learner matches and sometimes outperforms strong methods like ERM and GroupDRO. However, unlike these algorithms, it does not generalize well to other tasks. We show that it is possible to obtain an in-context learner that generalizes to unseen tasks by training on a diverse dataset of synthetic in-context learning instances.
@@ 1 Introduction
Large language models, such as GPT-3, have the ability of in-context learning (ICL), wherein they learn to solve a task given a few examples in the context (Brown et al., 2020). The most significant aspect of in-context learning is that the learning happens during the forward pass on the context and query, without updating network parameters. In order to study in-context learning in isolation, a number of studies considered training transformers (Vaswani et al., 2017) from scratch to solve simple learning tasks in-context. In particular, Garg et al. (2022) show empirically that transformers can be trained to perform in-context learning of simple regression functions, such as dense or sparse linear functions, two-layer ReLU neural networks, and small decision trees.
@@
Training on ICL instances can be seen as an instance of meta-learning (Schmidhuber, 1987; Naik and Mammone, 1992; Thrun and Pratt, 1998), where the goal is to learn a learning algorithm. What exact algorithm is learned when training transformers on ICL instances is still an open problem. Akyürek et al. (2022) and Von Oswald et al. (2023) show that transformers can implement a single gradient descent step of ordinary least squares and update the closed-form solution of ridge regression when a new example is added. Additionally, they provide evidence that transformers trained on ICL instances of linear regression learn algorithms that closely match predictions of the known algorithms, such as gradient descent on ordinary least squares objective and ridge regression. However, there is evidence that the learned algorithm may vary with model scale, depth, and pretraining task diversity (Akyürek et al., 2022; Raventós et al., 2024). In particular,  Raventós et al. (2024) demonstrate that in the setting of in-context learning of linear regression tasks with insufficient pretraining task diversity, the learned algorithm behaves like a Bayesian estimator with the pretraining task distribution as the prior, and hence fails to generalize well to unseen tasks. Yadlowsky et al. (2023) show that when trained on ICL instances where the regression function belongs to a union of distinct function classes, the learned algorithm fails to generalize beyond the pretraining function classes. Ahuja and Lopez-Paz (2023) show that in-context learning ability diminishes under strong distribution shifts.
@@
In this work, we explore the limits of in-context learning further by testing it on challenging settings. We deviate from the existing literature and consider visual classification tasks instead of regression tasks with simple function classes. In particular, we consider classification tasks where some features are spuriously correlated with the label. Such features are predictive of the label but are not causally related to it, due to which their correlation might not hold at test time. A prominent example is the cow vs camel classification task, where the background often correlates with the label, as cows are typically photographed in pastures, while camels are typically photographed in deserts (Beery et al., 2018). It is well-known that neural networks trained with gradient-based methods to minimize empirical risk can exploit spurious features, causing performance degradation under distribution shifts affecting these correlations (Torralba and Efros, 2011; Ribeiro et al., 2016; Gururangan et al., 2018; Zech et al., 2018; McCoy et al., 2019; Geirhos et al., 2019, 2020; Xiao et al., 2021).
@@
We start our analysis in the standard setting of having a single classification task with spurious features. We consider the conventional approach of obtaining an in-context learner, wherein a transformer is trained on sequences of form $(x_{1},y_{1},\ldots,x_{k},y_{k},x_{k+1})$ to predict the label $y_{k+1}$ of the query example $x_{k+1}$. We find that this conventional approach leads to models that do classification ignoring the context, essentially memorizing the task. Furthermore, these models lack robustness to changes of the correlation between the label and spurious features. In particular, we observe a significant performance drop when the query follows a distribution in which the label and spurious feature correlation is zero. We propose an effective approach of addressing the task memorization issue. Namely, we find that task memorization can be mitigated greatly by randomly permuting input embedding dimensions for each training sequence. To address the issue of spurious features, we propose a novel way of forming ICL instances and a suitable transformer architecture, which work together to simulate distribution shift with respect to spurious features in the context. Overall, our proposed techniques lead to strong in-context learners that outperform established methods such as 1-NN, empirical risk minimization (ERM), and GroupDRO (Sagawa* et al., 2020), suggesting that the in-context learner implements a more specialized algorithm.
@@
Despite being trained on instance of a single task, the learned algorithm generalizes to other tasks without spurious features. However, it fails to generalize to unseen tasks with spurious features. For this reason, we next explore training an in-context learner that generalizes to unseen tasks with spurious features. We create a dataset of in-context learning instances for various binary classifications tasks with varying spurious features. We demonstrate the efficacy of the proposed techniques on this dataset too and find that it can be improved further by passing spurious feature annotations as input and injecting occasional queries requesting the label of a proceeding context example to promote learning induction heads. The resulting model generalizes perfectly to unseen tasks, as long as the data generating process is similar. However, generalization to unseen tasks with possibly different data generating process depends on the severity of the challenge posed by spurious features, indicating that the learned algorithm is more brittle to severe distribution shifts than conventional algorithms. The source code for reproducing our experiments is available at https://github.com/YerevaNN/incontext_spurious.
@@
    We show that the conventional approach of training an in-context learner is susceptible to presence of spurious features and also leads to task memorization in case of a single task.
@@
    We propose a suite of novel techniques of forming in-context training data to mitigate task memorization and increase robustness to spurious features, leading to in-context learners that outperform established learning algorithms.
@@
(a) Naive approach
(b) Proposed approach
Figure 1: In-context learning transformer architectures of the naive and proposed approaches. The proposed approach allows arbitrary query tokens after each learning example. Token positions and attention mask are modified so that these intermediate queries have no effect on other tokens.
@@ 2 In-context learning based on a single task
We start by considering the common setting of having a single classification task with spurious features. For simplicity, we focus on label-balanced binary classification tasks in presence of a single binary spurious feature, although what follows next applies to label-imbalanced multiclass settings as well. Let $\mathcal{D}_{\mathrm{train}}$ be a set of training examples for the task, where each example is a triplet $(x,s,y)$ of input $x\in\mathbb{R}^{d}$, spurious feature value $s\in\left\{0,1\right\}$, and label $y\in\left\{0,1\right\}$. Similarly, let $\mathcal{D}_{\mathrm{test}}$ be a set of test examples. Importantly, we do not make any assumptions on the data generating process, except that $x$ has some information about $s$ and $s$ is predictive of $y$ on the training set, but their correlation does not hold on the test set. For an example $(x,s,y)$, we define its group $g=2y+s$. In a binary classification task with a single binary spurious feature, there are four groups. Without loss of generality, we assume that for a majority of training examples we have that $y=s$. Hence we refer to groups 0 and 3 as majority groups, while referring to groups 1 and 2 as minority groups.
Training a transformer to perform linear regression in-context requires millions of ICL training instances, even for small dimensional cases. For example, Garg et al. (2022) use 32 million training instances for 20-dimensional inputs. We next consider ways of generating so many ICL instances from a single task.
@@ 2.1 A naive approach of constructing ICL instances
The standard approach of constructing an ICL instance is to sample a subset of $n+1$ examples $\left\{(x_{i},s_{i},y_{i})\right\}_{i=1}^{n+1}$ from $\mathcal{D}_{\mathrm{train}}$ and form a sequence $S=(x_{1},\tilde{y}_{1},x_{2},\tilde{y}_{2},\ldots,x_{n},\tilde{y_{n}},x_{n+1})$, where $\tilde{y}_{i}\in\mathbb{R}^{d}$ is a fixed random representation of either $y_{i}$ or $g_{i}$ (this distinction will be elaborated later). Then one trains a transformer $f_{\theta}:\cup_{k}\mathbb{R}^{k\times d}\rightarrow[0,1]$ to predict $y_{i}$ given $S_{i}\triangleq(x_{1},\tilde{y}_{1},\ldots,x_{i-1},\tilde{y}_{i-1},x_{i})$ (see Figure 1(a)), optimizing the following loss function:
@@
where $\mathrm{CE}(y,\widehat{y})=-y\log\widehat{y}-(1-y)\log(1-\widehat{y})$ is the binary cross-entropy loss. We explore two options of setting $\tilde{y}_{i}$. In the first option, we set $\tilde{y}_{i}$ to represent $y_{i}$ with a constant vector or its negative in $\mathbb{R}^{d}$. In this case we aim to obtain an in-context learner that is robust to spurious features without receiving spurious feature annotations as input. ERM is one such learner that minimizes average loss on training examples and does not require spurious feature annotations. In the second option, we set $\tilde{y}_{i}$ to represent $g_{i}$ as a sum of two constant vectors in $\mathbb{R}^{d}$, one representing the class and the other representing the spurious feature. In this case we aim to obtain an in-context learner that does robust classification with respect to a specified spurious feature. GroupDRO is one such learner that minimizes worst-group loss, therefore requiring spurious feature annotations at training time.
@@
Unfortunately, the simple approach of (1) has several issues. First, as the classification task is the same in all ICL instances, the model can ignore context examples and predict $y_{i}$ based solely on $x_{i}$, essentially memorizing the task. Second, as all $n+1$ examples of a sequence $S$ are sampled from the training set and the spurious correlation holds for all of them, there is nothing preventing usage of spurious features in making predictions. To confirm these two issues, we consider the Waterbirds dataset (Sagawa* et al., 2020), which is landbird vs waterbird image classification task where image background (sea or land) is correlated with the label in the training set (4,795 examples), but not in validation and test sets . A robust classifier should predict waterbird or landbird without relying on image background. To separate out the representation learning challenge, we represent images with a pretrained and frozen DINOv2 ViT-B/14 distilled (Oquab et al., 2023). This way each image is embedded in $\mathbb{R}^{768}$. While using powerful pretrained representations increases overall performance under distribution shifts (Radford et al., 2021; Mehta et al., 2022), we note that it does not eliminate the problem of spurious correlations. Representations obtained via large-scale self-supervised pretraining are likely rich enough to capture information about both the label and spurious feature. Furthermore, many works have indicated that the main contribution to the out-of-domain generalization error comes from the classification head (rather than the representation learning module) and called for designing better methods of training the classification head (Galstyan et al., 2022; Menon et al., 2021; Kirichenko et al., 2023; Izmailov et al., 2022; Shi et al., 2023).
Figure 2: Majority-group and worst-group test accuracies on Waterbirds as a function of context size for the naive and proposed approaches with or without permuting input dimensions. Shaded regions show standard deviation across 5 training runs.
@@
We train a causal decoder-only GPT-J transformer (Wang and Komatsuzaki, 2021) with 80M parameters on 2M in-context learning sequence with $n=512$ and $\tilde{y}_{i}$ representing labels, constructed from the training set of Waterbirds. We use balanced sampling of classes and set the minority group proportion to 10% within each class. We use the ADAM optimizer (Kingma and Ba, 2014) ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) with 32 batch size and no weight decay. The learning rate is selected from $\left\{3\cdot 10^{-5},6\cdot 10^{-5},10^{-4}\right\}$ based on average test performance over 5 runs. Concretely, we evaluate on 8192 sequences where the context part is $n$ training examples, while the query is a sampled from the test set with equal group distribution. Exact metric definitions and missing details are provided in Appendix A. Note that with 512 context length and 10% minority group ratio within each class, the expected value of the number of context examples from each of the 2 minority groups is about 25. For reference, the smallest minority-group has only 56 examples in the Waterbirds training set.
@@
Figure 2 plots majority-group and worst-group test accuracies as a function of context size $n$. We see that naive approach results in models that ignore context – worst-group accuracy with 512 context examples is essentially the same as with 2 examples (see the naive curve). This confirms the task memorization issue. Figure 2 also shows that majority-group test accuracy of the naive approach is considerably higher compared to worst-group accuracy confirming the non-robustness issue.
@@ 2.2 The proposed approach of constructing ICL instances
Figure 3: Majority-group and worst-group test accuracies on Waterbirds-severe as a function of context size for the naive and proposed approaches with or without permuting input dimensions. Shaded regions show standard deviation across 5 training runs.
@@
To address the task memorization issue, we propose to rotate image embeddings in each ICL instance independently, making it harder to memorize individual examples. We found that generating random rotation matrices on fly is computationally expensive and slows down training. We tried generating and storing 10K rotation matrices, but this resulted in less than 50M different training examples that were still possible to memorize to some extent. A more effective and efficient alternative is to apply random permutations to image embedding dimensions (for brevity, this technique is denoted with +P in figures and tables). We found this approach to be very effective in terms of inducing in-context learning (see naive + P in Figure 2). We also see that the difference between majority-group and worst-group accuracies decreases, although an approximately 5 p.p. gap remains.
@@
When training an ICL transformer, ideally, we would like to simulate the situation of making a test prediction based on a context of training examples. Importantly, we would like to simulate the case where test distribution has balanced groups (i.e., the spurious correlation does not hold). Given access to spurious feature annotations for the training set, we can simulate this scenario using only training examples. In particular, we can form ICL instances of form $(x_{1},\tilde{y}_{1},\ldots,x_{n},\tilde{y}_{n},x_{n+1})$, where the context examples ($x_{1},\ldots,x_{n}$) are sampled in a way that the spurious feature is correlated with the label, while the query $x_{n+1}$ is sampled to have a uniform group distribution. However, if we again optimize the loss of (1), for context lengths less than $n$, the network will be allowed to make predictions using the spurious feature, which is undesirable. Please refer to Figure 13 of Appendix B for evidence of this. Potential ways of addressing this issue is upweighting the final prediction loss in Equation 1 or upweighting predictions on minority examples. In our preliminary experiments we found the former approach ineffective. We did not experiment with the latter approach.
Instead, we propose a novel way of forming in-context learning instances and a modified transformer architecture that is suitable for such sequences. In particular, we form sequences of form $S=(x_{1},\tilde{y}_{1},q_{1},x_{2},\tilde{y}_{2},q_{2},\ldots,x_{n},\tilde{y}_{n},q_{n})$, where $(x_{i},\tilde{y}_{i})$ are context examples, while $q_{i}$ are queries, sampled with replacement from $\mathcal{D}_{\mathrm{train}}\setminus\{x_{1},\ldots,x_{n}\}$. Importantly, $q_{i}$ are sampled with a uniform group distribution. Redefining $S_{i}=(x_{1},\tilde{y}_{1},q_{1},\ldots,x_{i},\tilde{y}_{i},q_{i})$, we would like the prediction on $S_{i}$ to be the label of $q_{i}$. When making a prediction on $q_{i}$, we want $q_{j}\ (j<i)$ to have no effect. For this end we make two modifications. First, we modify the causal attention matrix to disallow attending to query tokens, unless a query token is attending to itself. Formally, if we enumerate tokens from $1$ to $3n$ and define $M_{i,j}$ to denote the attention mask for token $i$ attending to token $j$, then we set
$M_{i,j}=\begin{cases}0,&i<j,\\ 0,&i>j\text{\ \ and\ \ }j\equiv 0\mod 3,\\ 1,&\text{otherwise.}\end{cases}$ (2)
@@
Second, we use modified token positions for computing positional encodings, in order to discount intermediate query tokens. Namely, for the sequence $(x_{1},\tilde{y}_{1},q_{1},x_{2},\tilde{y}_{2},\ldots,x_{n},\tilde{y}_{n},q_{n})$, position indices are set to $(0,1,2,2,3,4,4,\ldots,2n-2,2n-1,2n)$. Formally, enumerating tokens from $1$ to $3n$, the position index of the $i$-th token is set to $2\left\lfloor\frac{i-1}{3}\right\rfloor+(i-1)\mod 3$. Please refer to Figure 1 for an illustration. Hereafter, we refer to this approach as simply “proposed approach”.

### v2

@@ Abstract
Large language models exhibit a remarkable capacity for in-context learning, where they learn to solve tasks given a few examples. Recent work has shown that transformers can be trained to perform simple regression tasks in-context. This work explores the possibility of training an in-context learner for classification tasks involving spurious features. We find that the conventional approach of training in-context learners is susceptible to spurious features. Moreover, when the meta-training dataset includes instances of only one task, the conventional approach leads to in-weights learning and fails to produce a model that leverages context for predictions. Based on these observations, we propose a novel technique to train such a learner for a given classification task. Remarkably, this in-context learner matches and sometimes outperforms strong methods like ERM and GroupDRO. However, unlike these algorithms, it does not generalize well to other tasks. We show that it is possible to obtain an in-context learner that generalizes to unseen tasks by training on a diverse dataset of synthetic in-context learning instances.
@@ 1 Introduction
Large language models, such as GPT-3, have the ability of in-context learning (ICL), wherein they learn to solve a task given a few examples in the context (Brown et al., 2020). The most significant aspect of in-context learning is that the learning occurs during the forward pass on the context and query, without updating network parameters. In order to study in-context learning in isolation, a number of studies considered training transformers (Vaswani et al., 2017) from scratch to solve simple learning tasks in-context. In particular, Garg et al. (2022) show empirically that transformers can be trained to perform in-context learning of simple regression functions, such as dense or sparse linear functions, two-layer ReLU neural networks, and small decision trees.
@@
Training on ICL instances can be seen as an instance of meta-learning (Schmidhuber, 1987; Naik & Mammone, 1992; Thrun & Pratt, 1998), where the goal is to learn a learning algorithm. What exact algorithm is learned when training transformers on ICL instances is an open problem. Akyürek et al. (2023) and Von Oswald et al. (2023) show that transformers can implement a single gradient descent step of ordinary least squares and update the closed-form solution of ridge regression when a new example is added. Additionally, they provide evidence that transformers trained on ICL instances of linear regression learn algorithms that closely match predictions of the known algorithms, such as gradient descent on the ordinary least squares objective and ridge regression. However, there is evidence that the learned algorithm may vary with model scale, depth, and pretraining task diversity (Akyürek et al., 2023; Raventós et al., 2024; Goddard et al., 2025). In particular,  Raventós et al. (2024) demonstrate that in the setting of in-context learning of linear regression tasks with insufficient pretraining task diversity, the learned algorithm behaves like a Bayesian estimator with the pretraining task distribution as the prior, and hence fails to generalize well to unseen tasks. Yadlowsky et al. (2023) show that when trained on ICL instances where the regression function belongs to a union of distinct function classes, the learned algorithm fails to generalize beyond the pretraining function classes. Ahuja & Lopez-Paz (2023) show that in-context learning ability diminishes under strong distribution shifts.
In this work, we explore the limits of in-context learning further by studying it in a more challenging setting. In particular, motivated by results indicating that in-context learning is susceptible to shortcuts and spurious correlations (Tang et al., 2023; Zhou et al., 2024; Song et al., 2024), we pose the following question:
Can we obtain an effective classification algorithm that is robust to spurious features by training an in-context learner on a suitable meta-training dataset, rather than designing the learning algorithm manually?
@@
To address this question in isolation, we consider visual classification tasks where some features are spuriously correlated with the label. Such features are predictive of the label but are not causally related to it, due to which their correlation might not hold at test time. A prominent example is the cow vs. camel classification task, where the background often correlates with the label, since cows are typically photographed in pastures, while camels are typically photographed in deserts (Beery et al., 2018). It is well known that neural networks trained with gradient-based methods to minimize empirical risk can exploit spurious features, causing performance degradation under distribution shifts affecting these correlations (Torralba & Efros, 2011; Ribeiro et al., 2016; Gururangan et al., 2018; Zech et al., 2018; McCoy et al., 2019; Geirhos et al., 2019; 2020; Xiao et al., 2021).
@@
We start our analysis in the standard setting of having a single classification task with spurious features. We consider the conventional approach of obtaining an in-context learner, wherein a transformer is trained on sequences of form $(x_{1},y_{1},\ldots,x_{k},y_{k},x_{k+1})$ to predict the label $y_{k+1}$ of the query example $x_{k+1}$. We find that this conventional approach leads to in-weights learning (Chan et al., 2022), wherein models perform classification ignoring the context, essentially memorizing the task. Furthermore, these models lack robustness to changes in the correlation between the label and spurious features. In particular, we observe a significant performance drop when the query follows a distribution in which the label and spurious feature correlation is zero. We propose an effective approach to reduce in-weights learning and improve in-context learning. Namely, we find that in-weights learning can be greatly mitigated by randomly permuting input embedding dimensions for each training sequence. To address the issue of spurious features, we propose a novel way of forming ICL instances and a suitable transformer architecture, which work together to simulate distribution shift with respect to spurious features in the context. Overall, our proposed techniques lead to strong in-context learners that outperform established methods such as 1-NN, empirical risk minimization (ERM), and GroupDRO (Sagawa* et al., 2020), suggesting that the in-context learner implements a more specialized algorithm.
(a) Naive approach
(b) Proposed approach
Figure 1: In-context learning transformer architectures of the naive and proposed approaches. The proposed approach allows arbitrary query tokens after each learning example. Token positions and the attention mask are modified so that these intermediate queries have no effect on other tokens.
@@
Despite being trained on instances of a single task, the learned algorithm generalizes to other tasks without spurious features. However, it fails to generalize to unseen tasks with spurious features. For this reason, we next explore training an in-context learner that generalizes to unseen tasks with spurious features. We create a dataset of in-context learning instances for various binary classification tasks with varying spurious features. We demonstrate the efficacy of the proposed techniques on this dataset too and find that it can be improved further by passing spurious feature annotations as input and injecting occasional queries requesting the label of a preceding context example to promote learning induction heads. The resulting model generalizes perfectly to unseen tasks, as long as the data generating process is similar. However, generalization to unseen tasks with possibly different data generating process depends on the severity of the challenge posed by spurious features, indicating that the learned algorithm is more brittle to severe distribution shifts than conventional algorithms. The source code for reproducing our experiments is available at https://github.com/YerevaNN/incontext_spurious.
@@
    We show that the conventional approach of training an in-context learner is susceptible to presence of spurious features and also leads to in-weights learning in case of a single task.
@@
    We propose a suite of novel techniques of forming in-context training data to reduce in-weights learning and increase robustness to spurious features, leading to in-context learners that outperform established learning algorithms.
@@ 2 In-context learning based on a single task
(a) Waterbirds
(b) Waterbirds-severe
Figure 2: Worst-group test accuracy on Waterbirds and Waterbirds-severe as a function of context size for the naive and proposed methods with or without permuting input dimensions. Shaded regions show standard deviation across 5 training runs.
We start by considering the common setting of having a single classification task with spurious features. For simplicity, we focus on label-balanced binary classification tasks in presence of a single binary spurious feature, although what follows next applies to label-imbalanced multiclass settings as well. Let $\mathcal{D}_{\mathrm{train}}$ be a set of training examples for the task, where each example is a triplet $(x,s,y)$ of input $x\in\mathbb{R}^{d}$, spurious feature value $s\in\left\{0,1\right\}$, and label $y\in\left\{0,1\right\}$. Similarly, let $\mathcal{D}_{\mathrm{test}}$ be a set of test examples. Importantly, we do not make any assumptions on the data generating process, except that $x$ has some information about $s$ and $s$ is predictive of $y$ on the training set, but their correlation does not hold on the test set. For an example $(x,s,y)$, we define its group $g=2y+s$. In a binary classification task with a single binary spurious feature, there are four groups. Without loss of generality, we assume that for majority of training examples we have $y=s$. Hence, we refer to groups 0 and 3 as majority groups, while referring to groups 1 and 2 as minority groups.
Training a transformer to perform linear regression in-context requires millions of ICL training instances, even for small dimensional cases. For example, Garg et al. (2022) use 32 million training instances for 20-dimensional inputs. Next, we consider ways of generating so many ICL instances from a single task.
@@ 2.1 A naive approach of constructing ICL instances
The standard approach to constructing an ICL instance is to sample a subset of $n+1$ examples $\left\{(x_{i},s_{i},y_{i})\right\}_{i=1}^{n+1}$ from $\mathcal{D}_{\mathrm{train}}$ and form a sequence $S=(x_{1},\tilde{y}_{1},x_{2},\tilde{y}_{2},\ldots,x_{n},\tilde{y_{n}},x_{n+1})$, where $\tilde{y}_{i}\in\mathbb{R}^{d}$ is a fixed random representation of either $y_{i}$ or $g_{i}$ (this distinction will be elaborated on later). Then one trains a transformer $f_{\theta}:\cup_{k}\mathbb{R}^{k\times d}\rightarrow[0,1]$ to predict $y_{i}$ given $S_{i}\triangleq(x_{1},\tilde{y}_{1},\ldots,x_{i-1},\tilde{y}_{i-1},x_{i})$ (see Figure˜1(a)), optimizing the following loss function:
@@
[…]
@@
Unfortunately, the simple approach of (1) has several issues. First, as the mapping from inputs to labels is the same for all ICL training instances, the model can do in-weights learning, wherein the model encodes this stable mapping in its weights and classifies queries without using context examples. In other words, the learned algorithm does no in-context learning and predict $y_{i}$ based solely on $x_{i}$, essentially memorizing the task. Second, as all $n+1$ examples of a sequence $S$ are sampled from the training set and the spurious correlation holds for all of them, there is nothing preventing usage of spurious features in making predictions. To confirm these two issues, we consider the Waterbirds dataset (Sagawa* et al., 2020), which is a landbird versus waterbird image classification task where image background (sea or land) is correlated with the label in the training set (4,795 examples), but not in the validation and test sets. A robust classifier should predict waterbird or landbird without relying on image background. To separate out the representation learning challenge, we represent images with a pretrained and frozen DINOv2 ViT-B/14 distilled (Oquab et al., 2023). This way each image is embedded in $\mathbb{R}^{768}$. While using powerful pretrained representations increases overall performance under distribution shifts (Radford et al., 2021; Mehta et al., 2022), we note that it does not eliminate the problem of spurious correlations. Representations obtained via large-scale self-supervised pretraining are likely rich enough to capture information about both the label and spurious feature. Furthermore, many works have indicated that the main contribution to the out-of-domain generalization error comes from the classification head (rather than the representation learning module) and called for designing better methods of training the classification head (Galstyan et al., 2022; Menon et al., 2021; Kirichenko et al., 2023; Izmailov et al., 2022; Shi et al., 2023).
@@
We train a causal decoder-only GPT-J transformer (Wang & Komatsuzaki, 2021) with 80M parameters on 2M in-context learning sequences with $n=512$ and $\tilde{y}_{i}$ representing labels, constructed from the training set of Waterbirds. We use balanced sampling of classes and set the minority group proportion to 10% within each class. We use the ADAM optimizer (Kingma & Ba, 2014) ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) with 32 batch size and no weight decay. The learning rate is selected from $\left\{3\cdot 10^{-5},6\cdot 10^{-5},10^{-4}\right\}$ based on average test performance over 5 runs. Concretely, we evaluate on 8192 sequences where the context part is $n$ training examples, while the query is a sampled from the test set with equal group distribution. Exact metric definitions and missing details are provided in Appendix˜A. Note that with 512 context length and 10% minority group ratio within each class, the expected value of the number of context examples from each of the 2 minority groups is about 25. For reference, the smallest minority group has only 56 examples in the Waterbirds training set.
@@
Figure˜2(a) plots worst-group test accuracy as a function of context size $n$. We see that the naive approach results in models that ignore context – worst-group accuracy with 512 context examples is essentially the same as with 2 examples (see the naive curve). This confirms the in-weights learning issue. Figure˜17(a) also shows that majority-group test accuracy of the naive approach is considerably higher compared to worst-group accuracy, confirming the non-robustness issue.
@@ 2.2 The proposed approach of constructing ICL instances
To prevent in-weights learning and induce in-context learning instead, potentially enabling generalization to other tasks, one approach is to increase the number of training tasks/mappings. We propose rotating input embeddings in each ICL instance independently as a simple approach of deriving many tasks from a single source task. When training with rotated input embeddings, every ICL sequences receives a different permutation and thus represents a different mapping from inputs to labels. This naturally encourages in-context learning, where the model uses context examples to infer the mapping on the fly, instead of encoding a single mapping. We found that generating random rotation matrices on-the-fly is computationally expensive and slows down training. We tried generating and storing 10K rotation matrices, but this resulted in fewer than 50M different training examples that were still possible to memorize to some extent. A more effective and efficient alternative is to apply random permutations to image embedding dimensions (for brevity, this technique is denoted by +P in figures and tables; please see Figure˜12 for an illustration of this technique). We found this approach to be very effective in inducing in-context learning (see naive + P in Figure˜2(a)). We also see that the difference between majority-group and worst-group accuracies decreases, although an approximately 5 p.p. gap remains.
@@
When training an ICL transformer, ideally, we would like to simulate the situation of making a test prediction based on a context of training examples. Importantly, we would like to simulate the case where the test distribution has balanced groups (i.e., the spurious correlation does not hold). Given access to spurious feature annotations for the training set, we can simulate this scenario using only training examples. In particular, we can form ICL instances of form $(x_{1},\tilde{y}_{1},\ldots,x_{n},\tilde{y}_{n},x_{n+1})$, where the context examples ($x_{1},\ldots,x_{n}$) are sampled in such a way that the spurious feature is correlated with the label, while the query $x_{n+1}$ is sampled to have a uniform group distribution. However, if we again optimize the loss of (1), for context lengths less than $n$, the network will be allowed to make predictions using the spurious feature, which is undesirable. Please refer to Figure˜19 of Appendix˜C for evidence of this. To address this, one can compute loss only on the query token, ignoring all intermediate predictions. Unfortunately, this approach leads to reduced sample efficiency and slower learning.
Instead, we introduce a notion of intermediate queries, which allows us to simulate making predictions on test examples sampled from a balanced group distribution at all context lengths at once. Given training examples $\{(x_{i},\tilde{y}_{i})\}_{i=1}^{n}$, we sample test examples $\{q_{i}\}_{i=1}^{n}$ from $\mathcal{D}_{\mathrm{train}}\setminus\{x_{1},\ldots,x_{n}\}$ with uniform group distribution. We name these test examples $\{q_{i}\}_{i=1}^{n}$ intermediate queries. Then, we form a sequence $S=(x_{1},\tilde{y}_{1},q_{1},x_{2},\tilde{y}_{2},q_{2},\ldots,x_{n},\tilde{y}_{n},q_{n})$. Our goal is to do a single forward pass on this sequence $S$ and obtain the predictions of form $\hat{y}_{i}\triangleq f_{\theta}((x_{1},\tilde{y}_{1},x_{2},\tilde{y}_{2},\ldots,x_{i},\tilde{y}_{i},q_{i}))$ for all $i\in[n]$. To this end we make two modifications, essentially making representations of context tokens $x_{1},\tilde{y_{1}},\ldots,x_{n},\tilde{y}_{n}$ agnostic to the presence of the intermediate queries $q_{1},\ldots,q_{n}$. First, we modify the causal attention matrix to disable attention to query tokens, unless a query token is attending to itself. See Figure˜11 for an illustration for $n=3$. Formally, if we enumerate tokens from $1$ to $3n$ and define $M_{i,j}$ as the attention mask for token $i$ attending to token $j$, then we set
$M_{i,j}=\begin{cases}0,&i<j,\\ 0,&i>j\text{\ \ and\ \ }j\equiv 0\hskip-5.0pt\mod 3,\\ 1,&\text{otherwise.}\end{cases}$ (2)
@@
Second, we use modified token positions for computing positional encodings, in order to discount intermediate query tokens. Namely, for the sequence $(x_{1},\tilde{y}_{1},q_{1},x_{2},\tilde{y}_{2},\ldots,x_{n},\tilde{y}_{n},q_{n})$, position indices are set to $(0,1,2,2,3,4,4,\ldots,2n-2,2n-1,2n)$. Formally, enumerating tokens from $1$ to $3n$, the position index of the $i$-th token is set to $2\left\lfloor\frac{i-1}{3}\right\rfloor+(i-1)\hskip-5.0pt\mod 3$. Please refer to Figure˜1 for an illustration.
Finally, once the predictions on the intermediate queries are obtained, we optimize the average loss on the query examples:
$\frac{1}{n}\sum_{i=1}^{n}\mathrm{CE}(y_{q_{i}},\hat{y}_{i}),$ (3)
where $y_{q_{i}}$ is the label of query $q_{i}$. Hereafter, we refer to this approach as simply “proposed approach”.